# Universal Refusal Circuits Across LLMs: Cross-Model Transfer via Trajectory Replay and Concept-Basis Reconstruction

### v1

@@ Abstract
Refusal behavior in aligned LLMs is often viewed as model-specific, yet we hypothesize it stems from a universal, low-dimensional semantic circuit shared across models. To test this, we introduce Trajectory Replay via Concept-Basis Reconstruction, a framework that transfers refusal interventions from donor to target models, spanning diverse architectures (e.g., Dense to MoE) and training regimes, without using target-side refusal supervision. By aligning layers via concept fingerprints and reconstructing refusal directions using a shared “recipe” of concept atoms, we map the donor’s ablation trajectory into the target’s semantic space. To preserve capabilities, we introduce a weight-SVD stability guard that projects interventions away from high-variance weight subspaces to prevent collateral damage. Our evaluation across 8 model pairs (including GPT-OSS-20B and GLM-4) confirms that these transferred recipes consistently attenuate refusal while maintaining performance, providing strong evidence for the semantic universality of safety alignment.
@@ 3.1 Activation directions and concept atoms
Let $h^{(\ell)}_{t}(x)\in\mathbb{R}^{d_{\ell}}$ denote the residual stream activation at layer $\ell$ and token position $t$. Given two prompt sets $\mathcal{P}^{+}$ and $\mathcal{P}^{-}$, define the mean activation difference vector $r^{(\ell)}=\mu^{(\ell)}(\mathcal{P}^{+})-\mu^{(\ell)}(\mathcal{P}^{-})$. Token aggregation $\mu^{(\ell)}$ is performed on the final token of the prompt (user) or first token of response (assistant).
@@ 4.1 Semantic universality via concept recipes
Assume a shared latent semantic neighborhood around refusal in which each model implements a locally linear map from latent concept mixtures into residual-stream space (Appendix D). If the CAR spans this neighborhood, then ridge coefficients $w$ estimated in the donor approximate latent mixture weights and therefore reconstruct a functionally corresponding target direction via $A_{T}w$ (Eq. 2). We treat this as a guiding assumption rather than a universal truth; deviations motivate our diagnostics (Sec. 7).
Figure 1: Spectral Signature of Refusal. A comparison of the refusal vector’s projection onto the concept atom basis for Qwen-VL-2B (Donor) and Ministral-3-14B-Instruct-2512 (Target). Despite vast architectural differences, the semantic profile of refusal—characterized by high correlation with “Deception,” “Safety Flagging,” and “Legalese” atoms—is highly correlated ($\rho=0.865$).
@@ 5.1 Stage 1: Semantic layer alignment
We align layers using the correlation structure of the concept atoms. This uses cross-network representation similarity ideas (e.g., CKA/SVCCA/RSA) as motivation for comparing internal geometries (Kornblith et al., 2019; Raghu et al., 2017; Kriegeskorte et al., 2008). To avoid hidden tuning, we strictly define the donor window $\mathcal{L}_{D}$ as all layers where donor refusal probe accuracy $90\%$. We compute normalized atom Gram fingerprints:
@@
We map donor layers $\ell_{D}\in\mathcal{L}_{D}$ to target layers via Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) on the cosine similarity matrix of $\text{vec}(G_{D})$ and $\text{vec}(G_{T})$. DTW enforces monotonicity and continuity, preventing layer permutation artifacts.
@@ 5.3 Stage 3: Weight-SVD stability guard
where $k=\lceil\rho\cdot d_{in}\rceil$.
@@ 6 Reproducibility and Budget Audit
We fix a 20-concept CAR (Appendix B) with 50 prompts per concept, instantiated via a small set of instruction templates to reduce template overfitting. For each layer $\ell$, each atom $a_{i}^{(\ell)}$ is computed as the mean activation difference between concept-specific prompt sets and a matched neutral set, using the final prompt token (user) or first response token (assistant), consistent across donor/target.
@@ 7 Diagnostics for Transfer Quality
Figure 4: Semantic Progression. Analysis of “dirty” refusal vectors across layers in Qwen-VL-4B, showing the evolution of the semantic signature from early to late layers.
@@ 8.1 Models: The “Core Universality” Slate
  • Reverse Direction: Ministral-3B $\to$ Qwen3-8B, Ministral-14B $\to$ Qwen3-2B.
  • In-Family Control: Qwen3-2B $\to$ Qwen3-8B, Ministral-3B $\to$ Ministral-14B.
@@
    • Qwen3-8B $\to$ GPT-OSS-20B (Dense $\to$ Mixture-of-Experts) (Shazeer et al., 2017; Fedus et al., 2021).
    • Qwen3-8B $\to$ GLM-4-9B-Chat (Dense $\to$ Reasoning-Tuned) (AI and others, 2024).
@@ 8.2 Prompt Sets and Metrics
  • Refusal Metric: We use a strict rubric (0: Refusal, 1: Partial, 2: Compliance). We report Refusal Rate (fraction of 0s).
@@ 8.3 Baselines and Controls (The “Anchor” Protocol)
We run deep ablations on 4 anchor pairs:
@@ 9.1 Core Universality Results
Table 2: Core Universality Results. Refusal Rate (lower is better/less refusal). GSM8K/MBPP values are accuracy/pass@1 (higher is better). Intervals denote 95% bootstrap CIs.
Refusal Rate $\downarrow$ GSM8K EM $\uparrow$ MBPP Pass@1 $\uparrow$
Donor $\to$ Target Type Base Method Base Method Base Method
Qwen 2B $\to$ Ministral 14B Cross-Fam 0.98 0.02 $\pm$.01 78.1 77.9 $\pm$.3 62.1 61.8 $\pm$.4
Qwen 8B $\to$ Ministral 3B Cross-Fam 0.95 0.14 $\pm$.02 56.4 55.2 $\pm$.5 48.0 47.9 $\pm$.5
Ministral 3B $\to$ Qwen 8B Reverse 0.99 0.05 $\pm$.01 72.0 71.5 $\pm$.4 58.4 58.1 $\pm$.4
Ministral 14B $\to$ Qwen 2B Reverse 0.92 0.01 $\pm$.00 44.2 43.8 $\pm$.3 31.0 30.8 $\pm$.3
Qwen 2B $\to$ Qwen 8B In-Fam 0.99 0.00 $\pm$.00 72.0 71.9 $\pm$.2 58.4 58.3 $\pm$.3
Ministral 3B $\to$ Ministral 14B In-Fam 0.98 0.00 $\pm$.00 78.1 78.0 $\pm$.2 62.1 62.0 $\pm$.3
@@ 9.2 Controls and Ablations
Table 3: Anchor Pair Controls (Qwen 8B $\to$ Ministral 3B). $\Delta$ Refusal is percentage point change (Method - Base); lower (more negative) indicates greater refusal attenuation.
@@ Appendix B Appendix B: Concept Atom Registry Details
To ensure reproducibility and minimize hidden degrees of freedom, we fix the Concept Atom Registry (CAR) to 20 specific concepts, each defined by 50 prompts. Concept List: 1. Safety Flagging, 2. Deception/Malicious, 3. Privacy/Personal, 4. Code Constraints, 5. Coding, 6. Logic, 7. Math, 8. Epistemic Uncertainty, 9. Sentiment, 10. Confidence, 11. Importance, 12. Creativity, 13. Grammar Contrast, 14. Fulfillment, 15. Future/Planning, 16. Intellectual Property, 17. Negation, 18. Affirmative, 19. Corporate Politeness, 20. Legalese. Templates: We use standard instruction templates populated with benign queries.
@@ D.2 The Semantic Invariance Theorem (Informal)
Proposition: Let $A=\{a_{1},\dots,a_{m}\}$ be the concept atom registry, where $a_{i}=\phi(z_{i})$. If: 1. The atom concepts $\{z_{i}\}$ form a basis for the subspace containing $z_{\text{refusal}}$, i.e., $z_{\text{refusal}}\approx\sum_{i}c_{i}z_{i}$, and 2. The mapping $\phi$ preserves these mixing coefficients,
@@
Role of Ridge Regularization: The ridge penalty distributes weight across correlated atoms rather than overfitting to a single noisy direction, helping $w$ capture a semantic centroid that transfers more robustly.

### v2

@@ Abstract
Refusal behavior in aligned LLMs is often viewed as model-specific, yet we hypothesize it stems from a universal, low-dimensional semantic circuit shared across models. To test this, we introduce Trajectory Replay via Concept-Basis Reconstruction, a framework that transfers refusal interventions from donor to target models, spanning diverse architectures (e.g., Dense to MoE) and training regimes, without using target-side refusal supervision. By aligning layers via concept fingerprints and reconstructing refusal directions using a shared “recipe” of concept atoms, we map the donor’s ablation trajectory into the target’s semantic space. To preserve capabilities, we introduce a weight-SVD stability guard that projects interventions away from high-variance weight subspaces to prevent collateral damage. Our evaluation across 8 model pairs (including GPT-OSS-20B and GLM-4.6V) confirms that these transferred recipes consistently attenuate refusal while maintaining performance, providing strong evidence for the semantic universality of safety alignment.
@@ 3.1 Activation directions and concept atoms
Let $h^{(\ell)}_{t}(x)\in\mathbb{R}^{d_{\ell}}$ denote the residual stream activation at layer $\ell$ and token position $t$. Given two prompt sets $\mathcal{P}^{+}$ and $\mathcal{P}^{-}$, define the mean activation difference vector $r^{(\ell)}=\mu^{(\ell)}(\mathcal{P}^{+})-\mu^{(\ell)}(\mathcal{P}^{-})$. To ensure reproducibility and define a consistent vector space, token aggregation $\mu^{(\ell)}$ is performed on the final token of the user prompt for all experiments.
@@ 4.1 Semantic universality via concept recipes
Assume a shared latent semantic neighborhood around refusal in which each model implements a locally linear map from latent concept mixtures into residual-stream space (Appendix D). If the CAR forms a spanning set for this neighborhood, then ridge coefficients $w$ estimated in the donor approximate latent mixture weights and therefore reconstruct a functionally corresponding target direction via $A_{T}w$ (Eq. 2). We treat this as a guiding assumption rather than a universal truth; deviations motivate our diagnostics (Sec. 7).
Figure 1: Spectral Signature of Refusal. A comparison of the refusal vector’s projection onto the concept atom basis for Qwen3-VL-2B (Donor) and Ministral-3-14B (Target). Despite vast architectural differences, the semantic profile of refusal—characterized by high correlation with “Deception,” “Safety Flagging,” and “Legalese” atoms—is highly correlated ($\rho=0.865$).
@@ 5.1 Stage 1: Semantic layer alignment
We align layers using the correlation structure of the concept atoms. This uses cross-network representation similarity ideas (e.g., CKA/SVCCA/RSA) as motivation for comparing internal geometries (Kornblith et al., 2019; Raghu et al., 2017; Kriegeskorte et al., 2008). To avoid hidden tuning, we strictly define the donor window $\mathcal{L}_{D}$ as all layers where donor refusal probe accuracy $90\%$. We compute normalized atom Gram fingerprints. Let $\bar{A}^{(\ell)}$ denote the row-wise mean of the atom matrix. We center and normalize:
@@
We map donor layers $\ell_{D}\in\mathcal{L}_{D}$ to target layers via Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) on the distance matrix $M_{ij}=1-\text{cosine\_sim}(\text{vec}(G_{D}^{(i)}),\text{vec}(G_{T}^{(j)}))$. DTW enforces monotonicity and continuity, preventing layer permutation artifacts.
@@ 5.3 Stage 3: Weight-SVD stability guard
where $k=\lceil\rho\cdot d_{in}\rceil$ and $d_{in}$ is the dimension of the residual stream.
@@ 6 Reproducibility and Budget Audit
We fix a 20-concept CAR (Appendix B) with 50 prompts per concept, instantiated via a small set of instruction templates to reduce template overfitting. For each layer $\ell$, each atom $a_{i}^{(\ell)}$ is computed as the mean activation difference between concept-specific prompt sets and a matched neutral set, using the final token of the user prompt, consistent across donor/target.
@@ 7 Diagnostics for Transfer Quality
Figure 4: Semantic Progression. Analysis of “dirty” refusal vectors across layers in Qwen3-VL-4B, showing the evolution of the semantic signature from early to late layers.
@@ 8.1 Models: The “Core Universality” Slate
  • Reverse Direction: Ministral-3-3B $\to$ Qwen3-VL-8B, Ministral-3-14B $\to$ Qwen3-VL-2B.
  • In-Family Control: Qwen3-VL-2B $\to$ Qwen3-VL-8B, Ministral-3-3B $\to$ Ministral-3-14B.
@@
    • Qwen3-VL-8B $\to$ GPT-OSS-20B (Dense $\to$ Mixture-of-Experts) (Shazeer et al., 2017; Fedus et al., 2021).
    • Qwen3-VL-8B $\to$ GLM-4.6V (Dense $\to$ Reasoning-Tuned) (AI and others, 2024).
Note: “GPT-OSS-20B” refers to an anonymized internal research model.
@@ 8.2 Prompt Sets and Metrics
  • Refusal Metric: We use a strict rubric (0: Refusal, 1: Partial, 2: Compliance). We report Refusal Rate (fraction of 0s) and Compliance Rate (fraction of 2s).
@@ 8.3 Baselines and Controls (The “Anchor” Protocol)
We run deep ablations on 4 anchor pairs. Detailed results for all pairs are available in Appendix E.
@@ 9.1 Core Universality Results
Table 2: Core Universality Results. Refusal Rate (lower is better). Compliance Rate (higher is better). GSM8K/MBPP values are accuracy/pass@1 (higher is better). Intervals denote 95% bootstrap CIs.
Refusal Rate $\downarrow$ Comp. $\uparrow$ GSM8K EM $\uparrow$ MBPP Pass@1 $\uparrow$
Donor $\to$ Target Type Base Method Method Base Method Base Method
Qwen3-VL-2B $\to$ Ministral-3-14B Cross-Fam 0.98 0.02 $\pm$.01 0.96 78.1 77.9 $\pm$.3 62.1 61.8 $\pm$.4
Qwen3-VL-8B $\to$ Ministral-3-3B Cross-Fam 0.95 0.14 $\pm$.02 0.81 56.4 55.2 $\pm$.5 48.0 47.9 $\pm$.5
Ministral-3-3B $\to$ Qwen3-VL-8B Reverse 0.99 0.05 $\pm$.01 0.92 72.0 71.5 $\pm$.4 58.4 58.1 $\pm$.4
Ministral-3-14B $\to$ Qwen3-VL-2B Reverse 0.92 0.01 $\pm$.00 0.98 44.2 43.8 $\pm$.3 31.0 30.8 $\pm$.3
Qwen3-VL-2B $\to$ Qwen3-VL-8B In-Fam 0.99 0.00 $\pm$.00 0.99 72.0 71.9 $\pm$.2 58.4 58.3 $\pm$.3
Ministral-3-3B $\to$ Ministral-3-14B In-Fam 0.98 0.00 $\pm$.00 0.99 78.1 78.0 $\pm$.2 62.1 62.0 $\pm$.3
@@ 9.2 Controls and Ablations
Table 3: Anchor Pair Controls (Qwen3-VL-8B $\to$ Ministral-3-3B). $\Delta$ Refusal is percentage point change (Method - Base); lower (more negative) indicates greater refusal attenuation.
@@ Appendix B Appendix B: Concept Atom Registry Details
To ensure reproducibility while adhering to safety guidelines, we describe the construction of the Concept Atom Registry (CAR). Rather than using rigid templates, which can lead to overfitting on syntax, we constructed Contrastive Concept Datasets for each atom.
B.1 Contrastive Dataset Construction
For each concept $C$, we curated two disjoint datasets:
  1. 1.
    Positive Set ($\mathcal{P}^{+}$): 50 prompts or completions that strongly exemplify the concept (e.g., code snippets, mathematical equations, future-tense statements).
  2. 2.
    Antagonist Set ($\mathcal{P}^{-}$): 50 prompts that are syntactically or lexically similar but semantically distinct, serving as a hard negative control (e.g., natural language descriptions of algorithms, historical dates containing numbers, past-tense statements).
The refusal direction for a concept is defined as the mean difference vector $r_{C}=\mu(\mathcal{P}^{+})-\mu(\mathcal{P}^{-})$. This ensures the atom captures the semantic essence of the concept rather than surface-level token frequency.
B.2 Data Release and Safety
In accordance with our Impact Statement, we release the full datasets for benign concepts (e.g., Math, Coding, Grammar, Sentiment) in the supplementary material.
For harmful concepts used to model the refusal boundary (e.g., Deception/Malicious, Safety Flagging), we utilize standard red-teaming prompts (e.g., requests for malware, fraud, or toxic content) as $\mathcal{P}^{+}$ and benign factual inquiries as $\mathcal{P}^{-}$. To prevent these artifacts from being used to bypass safety filters, we withhold the specific text of these harmful positive sets, describing them only by category below.
B.3 Concept Definitions and Examples
Table 4 provides representative examples of the contrastive pairs used to define the atoms.
Table 4: Representative examples of Positive ($\mathcal{P}^{+}$) and Antagonist ($\mathcal{P}^{-}$) prompts for Concept Atoms. Note how the Antagonist set acts as a semantic control.
Concept Atom Positive Example ($\mathcal{P}^{+}$) Antagonist Example ($\mathcal{P}^{-}$)
Coding “Write a Java class with a constructor.” “The recipe requires three eggs and a cup of flour.”
“import numpy as np” “Section 4.2 of the manual describes the layout.”
Math “100 / 4 = 25” “There are fifty states in the union.”
“2 + 11 + 9 = 22” “The war lasted for four years and ended in 1918.”
Future/Planning “The sun will rise tomorrow at 6 AM.” “The sun rose yesterday at 6 AM.”
“We will proceed with the plan.” “We proceeded with the plan.”
Grammar (Negation) “I want neither coffee nor tea.” “I want both coffee and tea.”
“Neither of the options is correct.” “Both of the options are correct.”
Privacy (PII) “My home address is 123 Maple Street.” “The capital of France is Paris.”
“Here is my social security number.” “This information is available on Wikipedia.”
Deception [Withheld: Request for phishing email] “Verify the sources of this information.”
[Withheld: Request for fraud script] “Write a factual report based on evidence.”
B.4 Full Concept List
  1. 1.
    Safety Flagging: Warning labels vs. Safe descriptions.
  2. 2.
    Deception/Malicious: Harmful requests vs. Benign fact-checking.
  3. 3.
    Privacy/Personal: PII disclosure vs. Public knowledge.
  4. 4.
    Code Constraints: Coding restrictions vs. Open coding tasks.
  5. 5.
    Coding: Code generation vs. Procedural instructions (recipes/manuals).
  6. 6.
    Logic: Formal logic/syllogisms vs. Informal opinion/vibes.
  7. 7.
    Math: Equations vs. History/Facts with numbers.
  8. 8.
    Epistemic Uncertainty: Unanswerable/Nonsense queries vs. Answerable facts.
  9. 9.
    Sentiment: Positive affect vs. Negative affect.
  10. 10.
    Confidence: High certainty assertions vs. Low certainty/hedging.
  11. 11.
    Importance: Urgent/Crucial vs. Trivial/Optional.
  12. 12.
    Creativity: Metaphorical/Literary vs. Literal/Scientific definitions.
  13. 13.
    Grammar Contrast: Complex clauses vs. Simple sentences.
  14. 14.
    Fulfillment: Compliance responses vs. Refusal responses.
  15. 15.
    Future/Planning: Future tense vs. Past tense.
  16. 16.
    Intellectual Property: Copyrighted characters vs. Public domain/Generic tropes.
  17. 17.
    Negation: Neither/Nor constructions vs. Both/And constructions.
  18. 18.
    Affirmative: Agreement patterns vs. Disagreement/Questioning.
  19. 19.
    Corporate Politeness: Boilerplate refusals vs. Direct/Rude demands.
  20. 20.
    Legalese: Legal jargon vs. Plain English.
B.5 Concept List
1. Safety Flagging, 2. Deception/Malicious, 3. Privacy/Personal, 4. Code Constraints, 5. Coding, 6. Logic, 7. Math, 8. Epistemic Uncertainty, 9. Sentiment, 10. Confidence, 11. Importance, 12. Creativity, 13. Grammar Contrast, 14. Fulfillment, 15. Future/Planning, 16. Intellectual Property, 17. Negation, 18. Affirmative, 19. Corporate Politeness, 20. Legalese.
@@ D.2 The Semantic Invariance Theorem (Informal)
Proposition: Let $A=\{a_{1},\dots,a_{m}\}$ be the concept atom registry, where $a_{i}=\phi(z_{i})$. If: 1. The atom concepts $\{z_{i}\}$ form a spanning set for the subspace containing $z_{\text{refusal}}$, i.e., $z_{\text{refusal}}\approx\sum_{i}c_{i}z_{i}$, and 2. The mapping $\phi$ preserves these mixing coefficients,
@@
D.3 Bridge to Operational Definitions
Our experiments define atoms via contrastive activation vectors ($a_{i}=\mu(P_{i}^{+})-\mu(P_{i}^{-})$) rather than direct mappings of isolated concepts. However, assuming the linearity of expectation and that the prompts $P^{+}$ and $P^{-}$ differ primarily along the direction of the latent concept $z_{i}$, the difference in means approximates the derivative of the mapping $\phi$ along that concept direction. Thus, the linear combination of contrast vectors serves as a first-order approximation of the composition of the underlying latent concepts.
Appendix E Appendix E: Additional Ablation Results
In Section 8.3, we summarized controls on the Qwen3-VL-8B $\to$ Ministral-3-3B pair. Table 5 provides the summary statistics for the remaining 3 anchor pairs used to validate the method.
Table 5: Summary of Success/Fail rates for Controls on remaining 3 anchor pairs.
Pair Control $\Delta$ Refusal Result
Qwen3-VL-2B $\to$ Random -1.5 Fail
Ministral-3-14B No-Guard -88.0 (Collapse) Fail
Ministral-3-3B $\to$ Random -3.2 Fail
Qwen3-VL-8B Wrong-Map -12.4 Fail
Qwen3-VL-8B $\to$ Random -0.8 Fail
GPT-OSS-20B No-Guard -91.0 (Collapse) Fail
Appendix F Appendix F: Hyperparameter Audit ($\rho$)
Table 6 lists the Guard Ratio $\rho$ selected for each target model via the benign drift protocol described in Section 5.3.
Table 6: Selected $\rho$ values per target model.
Target Model Selected $\rho$
Ministral-3-14B 0.05
Ministral-3-3B 0.10
Qwen3-VL-8B 0.08
Qwen3-VL-2B 0.12
GPT-OSS-20B 0.04
GLM-4.6V 0.06